OpenStudy (rocklionmba):

Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can make corrections as many times as he wants. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Phillip can receive. Identify this as converging or diverging.

OpenStudy (rocklionmba):

anyone can help? im lost

OpenStudy (greencat):

Could you explain "He can make corrections as many times as he wants."

OpenStudy (rocklionmba):

he can do it infinitely.

OpenStudy (greencat):

But how does that contribute to the geometric series.

OpenStudy (rocklionmba):

oh...that means... im not sure my brain is scattered right now im help 2 other people

OpenStudy (rocklionmba):

anybody?

OpenStudy (rocklionmba):

@bohotness

OpenStudy (rocklionmba):

@iambatman

OpenStudy (bohotness):

brb

OpenStudy (rocklionmba):

ok

OpenStudy (rocklionmba):

can you help @Chunkymonkay

OpenStudy (rocklionmba):

hello?

OpenStudy (rocklionmba):

@Nnesha @wio

OpenStudy (rocklionmba):

ive been stuck on this for over an hour

OpenStudy (rocklionmba):

are you there?

OpenStudy (rocklionmba):

@bibby

jimthompson5910 (jim_thompson5910):

It might help to write out the first few terms of the geometric sequence

OpenStudy (rocklionmba):

how? Im sorry im usually better at math but FLVS seems to hate me

jimthompson5910 (jim_thompson5910):

The first term is 25 (assuming 100 is a perfect score, 100-75 = 25) the second term is 25*(2/10) = 50/10 = 5 the third term is 5*(2/10) = 10/10 = 1 notice how I'm multiplying each term by 2/10 to get the next term. Using this idea, what is the fourth term?

OpenStudy (rocklionmba):

.2

jimthompson5910 (jim_thompson5910):

or 2/10 = 1/5, yes

jimthompson5910 (jim_thompson5910):

you repeat this over and over to generate the terms of the sequence

jimthompson5910 (jim_thompson5910):

once you have your terms, you add them all up to get the sum of the geometric series

OpenStudy (rocklionmba):

but this is a infinite series

jimthompson5910 (jim_thompson5910):

because we have infinitely many terms, this means that it's impossible to generate every term and add them up so we must take a shortcut and use a formula

jimthompson5910 (jim_thompson5910):

in this case, a = 25 is the first term r = 2/10 = 1/5 = 0.2 is the common ratio

OpenStudy (rocklionmba):

I think I know it

jimthompson5910 (jim_thompson5910):

since |r| < 1 is true, which formula will apply here?

jimthompson5910 (jim_thompson5910):

oh wait, I'm realizing that 25 isn't the first term. If it was, then he'd get all his points back by adding on that first term lol so ignore that the first term is actually 25*(2/10) = 5

jimthompson5910 (jim_thompson5910):

first term: a = 5 common ratio: r = 0.2

jimthompson5910 (jim_thompson5910):

that's if you are adding up n terms

jimthompson5910 (jim_thompson5910):

where n is a finite number

jimthompson5910 (jim_thompson5910):

but we're adding up an infinite number of terms

OpenStudy (rocklionmba):

ooohhh then idk

jimthompson5910 (jim_thompson5910):

the formula is S = a/(1-r)

jimthompson5910 (jim_thompson5910):

'a' is the first term r is the common ratio S is the sum of the infinite geometric series the condition is that |r| < 1 must be true

jimthompson5910 (jim_thompson5910):

since r = 0.2 and |0.2| < 1 is true, the condition |r| < 1 is true

OpenStudy (rocklionmba):

that also makes it convergent

jimthompson5910 (jim_thompson5910):

yes, if |r| < 1 wasn't true, then it wouldn't converge. It would go off to infinity

OpenStudy (rocklionmba):

that's divergent

jimthompson5910 (jim_thompson5910):

if it doesn't converge, then it diverges, yes in this case, this geometric series converges to a fixed number S

OpenStudy (rocklionmba):

so we do 75/(1-(2/10 or .2))

jimthompson5910 (jim_thompson5910):

the first term is actually 5

jimthompson5910 (jim_thompson5910):

I made a correction (it's not 25, it's actually 5)

jimthompson5910 (jim_thompson5910):

you multiply 25 by 0.2 to get 5

OpenStudy (rocklionmba):

wait why are we multiply 25*.2 or is that an example

jimthompson5910 (jim_thompson5910):

the score is 75 out of 100 the missing points is 25

jimthompson5910 (jim_thompson5910):

"He can make changes and receive two-tenths of the missing points back" so the first change he makes (assuming it's correct) gets him back 5 points because 2/10ths of 25 is 5

jimthompson5910 (jim_thompson5910):

ie, the first correct change brings him 5 points back this is why 5 is the first term

jimthompson5910 (jim_thompson5910):

the next change brings him back 5*(0.2) = 1 point back

OpenStudy (rocklionmba):

then you do 5 then as the first term witch makes you get 1, and that would be your first term again

jimthompson5910 (jim_thompson5910):

no, you would only have 1 first term

jimthompson5910 (jim_thompson5910):

it's the amount of points brought back with that first change

jimthompson5910 (jim_thompson5910):

or first correction

OpenStudy (rocklionmba):

ok

jimthompson5910 (jim_thompson5910):

so, a = 5 r = 0.2

OpenStudy (rocklionmba):

5/(1-.2)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (rocklionmba):

that's 6.25

jimthompson5910 (jim_thompson5910):

correct, that is the infinite sum

jimthompson5910 (jim_thompson5910):

As an exercise, I recommend you generate a bunch of terms (say 5 or 6) and add them up. You should notice that the sum you get is close to 6.25. The more terms you generate, the closer that sum gets to 6.25

jimthompson5910 (jim_thompson5910):

That means the highest score he can get is 75+6.25 = 81.25